OpenStudy (andiemarinn):

Let f(x) = - √x+ 7. The y-intercept of f(x) is _____ and the zero is _____. A. (√7, 0); x =-7 B. (0, 7); x = -49 C. (0, 7); x = 49 D. (0, -7); x = 49 E. (0, 0); x = -√7

OpenStudy (legomyego180):

First to find the x-intercept set y=0

OpenStudy (andiemarinn):

So how would you plug in y?

OpenStudy (legomyego180):

\[0=-\sqrt{x}+7\] solve for x to get the intercept(s)

OpenStudy (legomyego180):

\[-7=-\sqrt{x}\] \[7=\sqrt{x}\] \[49 = x\] Right?

OpenStudy (legomyego180):

0-7=-7

OpenStudy (andiemarinn):

That's not right thank u

OpenStudy (legomyego180):

The y-intercept is 49